OpenStudy (anonymous):

Find the axis of symmetry for y = 4x^2 + 16x - 2

OpenStudy (anonymous):

do you know to do complete square method ?

OpenStudy (anonymous):

test -1 (-b/2a) = -16/2*4=-1

OpenStudy (paki):

The axis of symmetry would be found by -b/ 2a... have a try now please,,,,

Parth (parthkohli):

The axis of symmetry is just the vertical line passing through the vertex. In other words, since the vertex has coordinates \(\left(-b/2a, f(-b/2a)\right)\), you have the axis of symmetry as the line \(x = -b/2a\). Don't forget to include "\(x = \)"

OpenStudy (anonymous):

Do I use the formula ax^2 + bx + c =0

OpenStudy (anonymous):

Is the answer -8?

OpenStudy (anonymous):

Is it really?

OpenStudy (anonymous):

nope -18

OpenStudy (anonymous):

you could just use the equation givenby @ParthKohli

OpenStudy (anonymous):

I did.

OpenStudy (anonymous):

or go for complete square method, which i assume is not necassary if you can do it a simple way

OpenStudy (anonymous):

show me your working

OpenStudy (anonymous):

16x/2

OpenStudy (anonymous):

x=-8

OpenStudy (anonymous):

x = -b/2a b = 16 a = 4

OpenStudy (anonymous):

UGH This ish is so frustrating.

OpenStudy (anonymous):

why dont u substitute for a and b and find x :/ not so hard as think